# AgriFM: A Multi-source Temporal Remote Sensing Foundation Model for Crop Mapping

### v2

@@ 8 Conclusion
This paper first addresses and validates the necessity of simultaneous hierarchical spatiotemporal feature extraction for crop mapping. It leads to the development of a modified Video Swin Transformer architecture where temporal down-sampling is synchronized with spatial down-sampling operations. Capitalizing on this finding, we introduce AgriFM, a multi-source temporal remote sensing foundation model with specialized spatiotemporal modeling for agricultural crop mapping. AgriFM integrates multi-source satellite data including MODIS, Landsat-8/9, and Sentinel-2, and employs geographical prior land cover products for supervised pre-training. It effectively addresses the limitations of existing RSFMs, ensuring the incorporation of multi-resolution data, providing comprehensive coverage of the crop growth cycle, optimizing the use of geographical information, and offering a unified framework for a wide array of crop mapping tasks. These tasks include, but are not limited to, cropland mapping, cropland boundary delineation, early-season crop mapping, and specific tasks like winter wheat and rice mapping. The effectiveness and adaptability of AgriFM have been substantiated through superior performance across a variety of crop mapping tasks, demonstrating its potential as an immediately applicable solution for leveraging deep learning in more precise and efficient crop mapping.
@@
Our results confirm that pretraining significantly boosts performance, especially for complex tasks. However, architectural suitability remains equally crucial. This study’s practical implication is clear: for crop mapping tasks, Video Swin Transformer should be the architecture of first consideration, with domain-specific pretraining then applied for additional gains. This two-stage approach—selecting the optimal architecture followed by targeted pretraining—proves more effective than relying solely on either component.
@@
Codes will be available on the website of the Jockey Club STEM Lab of Quantitative Remote Sensing, HKU (https://glass.hku.hk) and https://github.com/flyakon/AgriFM.

### v3

@@ 8 Conclusion
This paper first addresses and validates the necessity of simultaneous hierarchical spatiotemporal feature extraction for crop mapping. It leads to the development of a modified Video Swin Transformer architecture where temporal down-sampling is synchronized with spatial down-sampling operations. Capitalizing on this finding, we introduce AgriFM, a multi-source temporal remote sensing foundation model with specialized spatiotemporal modeling for agricultural crop mapping. AgriFM integrates multi-source satellite data including MODIS, Landsat-8/9, and Sentinel-2, and employs geographical land cover products for supervised pre-training. It effectively addresses the limitations of existing RSFMs, ensuring the incorporation of multi-resolution data, providing comprehensive coverage of the crop growth cycle, optimizing the use of geographical information, and offering a unified framework for a wide array of crop mapping tasks. These tasks include, but are not limited to, agricultral land mapping, field boundary delineation, and specific tasks like winter wheat and rice mapping. The effectiveness and adaptability of AgriFM have been substantiated through superior performance across a variety of crop mapping tasks, demonstrating its potential as an immediately applicable solution for leveraging deep learning in more precise and efficient crop mapping.
@@
Our results confirm that pretraining significantly boosts performance, especially for complex tasks. However, architectural suitability remains equally crucial. This study’s practical implication is clear: for agriculture mapping tasks, Video Swin Transformer should be the architecture of first consideration, with domain-specific pretraining then applied for additional gains. This two-stage approach—selecting the optimal architecture followed by targeted pretraining—proves more effective than relying solely on either component.
@@
Codes and models are available on the website of the Jockey Club STEM Lab of Quantitative Remote Sensing, HKU (https://glass.hku.hk) and https://github.com/flyakon/AgriFM.